PLEASE SOLVE USING EXCELYou are in charge of loading cargo ships for ICC at a major east coast port. You have been asked to prepare a loading plan fo
Answer #1

a. Let,

C1F = number of commodity 1 to be loaded in forward position of Cargo ship

C1C = number of commodity 1 to be loaded in center position of Cargo ship

C1R = number of commodity 1 to be loaded in rear position of Cargo ship

C2F = number of commodity 2 to be loaded in forward position of Cargo ship

C2C = number of commodity 2 to be loaded in center position of Cargo ship

C2R = number of commodity 2 to be loaded in rear position of Cargo ship

C3F = number of commodity 3 to be loaded in forward position of Cargo ship

C3C = number of commodity 3 to be loaded in center position of Cargo ship

C3R = number of commodity 3 to be loaded in rear position of Cargo ship

C4F = number of commodity 4 to be loaded in forward position of Cargo ship

C4C = number of commodity 4 to be loaded in center position of Cargo ship

C4R = number of commodity 4 to be loaded in rear position of Cargo ship

Objective is to maximize profit so objective function = Max 70(C1F+C1C+C1R) + 50(C2F+C2C+C2R) + 60(C3F+C3C+C3R) + 80(C4F+C4C+C4R)

Subject to

C1F+C1C+C1R <= 3000 (3000 tons available to be shipped for commodity 1)

C2F+C2C+C2R <= 4000   (4000 tons available to be shipped for commodity 2)

C3F+C3C+C3R <= 1000 (1000 tons available to be shipped for commodity 3)

C4F+C4C+C4R <= 3000   (3000 tons available to be shipped for commodity 4)

C1F+C2F+C3F+C4F <= 3000 (weight capacity is 3000 tons in Forward hold)

C1C+C2C+C3C+C4C <= 5000 (weight capacity is 5000 tons in center hold)

C1R+C2R+C3R+C4R <= 2000 (weight capacity is 2000 tons in rear hold)

40C1F+25C2F+60C3F+50C4F <= 100000 (Volume capacity is 100000 in Forward hold)

40C1C+25C2C+60C3C+50C4C <= 160000 (Volume capacity is 160000 in center hold)

40C1R+25C2R+60C3R+50C4R <= 130000 (Volume capacity is 130000 in rear hold)

C1R+C2R+C3R+C4R >= 0.45(C1R+C2R+C3R+C4R+C1C+C2C+C3C+C4C+C1F+C2F+C3F+C4F)

C1R+C2R+C3R+C4R <= 0.55(C1R+C2R+C3R+C4R+C1C+C2C+C3C+C4C+C1F+C2F+C3F+C4F)

All variables greater than 0

Solving in excel solver we get,

Loading plan

C1F = number of commodity 1 to be loaded in forward position of Cargo ship = 0

C1C = number of commodity 1 to be loaded in center position of Cargo ship = 0

C1R = number of commodity 1 to be loaded in rear position of Cargo ship = 1444.444

C2F = number of commodity 2 to be loaded in forward position of Cargo ship = 0

C2C = number of commodity 2 to be loaded in center position of Cargo ship = 0

C2R = number of commodity 2 to be loaded in rear position of Cargo ship = 0

C3F = number of commodity 3 to be loaded in forward position of Cargo ship = 0

C3C = number of commodity 3 to be loaded in center position of Cargo ship = 0

C3R = number of commodity 3 to be loaded in rear position of Cargo ship = 0

C4F = number of commodity 4 to be loaded in forward position of Cargo ship = 1269.841

C4C = number of commodity 4 to be loaded in center position of Cargo ship = 1174.603

C4R = number of commodity 4 to be loaded in rear position of Cargo ship = 555.5556

and maximum profit = $341111.1

b. Sensitivity report

Variable Cells Final Value Reduced Objective Allovable Allov able Cost oefficien Increase Decrease 1E+30 1E+30 Cell Name $C$2

We can see from sensitivity report, only Volume capacity and Weight capacity of Rear hold have been fully utilized. Shadow price of Weight capacity of Rear hold is higher than volume capacity of rear hold. So expanding Weight capacity of rear hold would have maximized effect on Profit. 1 unit of Weight capacity increase of rear hold will fetch $155.55556 additional profit

Marginal value of Volume capacity - rear hold expansion is $1.721E-15

Marginal value of Weight capacity - rear hold expansion is $155.55556

Marginal value of other expansions is 0 as their existing capacities are not utilized